Inventory records for Marvin Company revealed the following: Date Transaction Number of Units Unit Cost Mar. 1 Beginning Inventory 1,000 $ 7.20 Mar. 10 Purchase 600 7.25 Mar. 16 Purchase 800 7.30 Mar. 23 Purchase 600 7.35 Marvin sold 2,300 units of inventory during the month. Ending inventory assuming LIFO would be: 2. Inventory records for Marvin Company revealed the following: Date Transaction Number of Units Unit Cost Mar. 1 Beginning inventory 990 $ 7.29 Mar. 10 Purchase 580 7.79 Mar. 16 Purchase 468 8.39 Mar. 23 Purchase 550 9.09 Marvin sold 1,910 units of inventory during the month. Ending inventory assuming FIFO would be: (Do not round your intermediate calculations. Round your answer to the nearest dollar amount.)m  
Answer #1

Under LIFO method, the ending inventory consists of the units from earliest purchases.

Units available for sale = 1,000 + 600 + 800 + 600 = 3,000

Ending inventory units = Units available for sale - Units sold

= 3,000 - 2,300

= 700

700 units in ending inventory consists of beginning inventory.

Ending inventory = 700 * $7.2 = $5,040

---------------------------------------------------------------------------------------------

Under FIFO method, the ending inventory consists of the units from recent purchases.

Units available for sale = 990 + 580 + 468 + 550 = 2,588

Ending inventory units = Units available for sale - Units sold

= 2,588 - 1,910

= 678

678 units in ending inventory consists of 550 units from March 23 purchases and 128 units from March 16 purchases

Ending inventory = (550*$9.09) + (128*$8.39)

= $6,073